# RiboGen: RNA Sequence and Structure Co-Generation with Equivariant MultiFlow

### v4

@@ 1 Introduction
Ribonucleic acid (RNA) is a fundamental biomolecule that stands at the intersection of modern biology and the origins of life. RNA has proven to be a versatile molecule, playing key roles in messaging (Crick, 1970), catalytic functions (Altman & Guerrier-Takada, 1983), regulation, and diverse biological processes through its complex 3D structures (Fire et al., 1998). While traditional computational methods faced limitations in decoding RNA structure and facilitating RNA design, deep learning emerged as powerful approach to accurately predict RNA structures, enhance RNA engineering, and unlock new insights into its functional roles. Existing deep learning models often predict RNA structure from sequence or design sequences for target structures independently. However, the ability to simultaneously generate both sequence and structure remains a largely unexplored space in deep learning-based modeling of RNA. This co-generation capability can enable exploration of the sequence-structure landscape in novel ways. To address this gap, in this paper we introduce RiboGen for joint generation of all-atom structure and sequence of RNA. RiboGen is based on a Multiflow (Campbell et al., 2024) model that draws on Flow Matching (Lipman et al., 2022; Liu et al., 2022) and Discrete Flow (Gat et al., 2024; Campbell et al., 2024) for its generation. We train a large model and evaluate its chemical validity and self-consistency. Our results showcase the capabilities of RiboGen for generation and highlight that co-generation models are a promising avenue for modeling of RNA.
@@ 1.1 Related Work
Previous work in deep learning for RNA design has made significant progress in computational approaches to predict RNA structures (Shen et al., 2024; Abramson et al., 2024), RNA-RNA interactions, and the design of novel RNA sequences. Recent progress in generative modeling for RNA focuses on sequence and structure generation through Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) or Flow Matching (Lipman et al., 2022). MMDiff (Morehead et al., 2023a) uses discrete DDPM to co-generate sequence and structure of RNA, DNA and proteins. Our approach instead employs Flow Matching and its discrete variant (Campbell et al., 2024). RNA-FrameFlow (Anand et al., 2024) represents RNA through rigid body frames and uses flow matching to generate 3D backbones, employing inverse folding model gRNAde (Joshi et al., 2023) to obtain sequences. In contrast, while similarly using Flow Matching for 3D generation, our approach additionally models the discrete sequence components of RNA generation. RNAFlow Nori & Jin (2024) uses a GNN conditioned on protein structure and sequence to generate RNA sequences, which are then processed by RoseTTAFold2NA Baek (2024) to predict backbone structure; this method additionally conditions on protein structure as input. Our approach instead focuses on isolated RNA, learning unconditional direct sequence-structure generation. Recent application of the Multiflow (Campbell et al., 2024) framework to protein sequence-structure design has demonstrated the power of joint generation. Our approach builds upon on these insights from protein design, and adapts them to the RNA design domain by enabling the joint generation of RNA sequences and all-atom structures.
@@ 2.2 Flow Matching
To model the distribution of 3D RNA coordinates and features we use Flow Matching (Lipman et al., 2022; Liu et al., 2022; Albergo et al., 2023). Flow Matching parameterizes a conditional probability path on time by learning a conditional velocity field that transforms samples from a prior distribution to a target data distribution . To learn this transport, we use the standard form of Flow Matching to obtain a noised version of via the linear interpolant and its associated velocity:
We build our model to reconstruct the target from its noised counterpart . We follow the reparameterization of (Jing et al., 2024; Pooladian et al., 2023) and obtain the learned conditional velocity through:
@@ 2.3 Discrete Flow Matching
While the standard form of Flow Matching is effective for continuous data, it is not appropriate for categorical domains. Hence, for modeling the RNA sequence we employ the extended framework of Discrete Flow Matching (Gat et al., 2024; Campbell et al., 2024). In this setting, the sequence data is described over a vocabulary . We parameterize a discrete flow by describing the velocity field over a probability vector on this categorical space:
@@ 2.4 Multiflow
To generate full RNA representations, we use Multiflow (Campbell et al., 2024) and train a neural network to learn the multimodal velocity field given jointly noised data and time (Figure 1.b). We employ the standard Flow Matching for and and its discrete counterpart for . This decomposition enables the model to separately capture the distributions of sequence, backbone, and atomic positions, allowing for unconditional generation or conditional generation based on specific structural or sequence constraints (Figure 2.a), such as structure prediction or inverse folding.
@@ 2.5 Architecture and Training
We use Euclidean-Equivariant Neural Networks (Geiger & Smidt, 2022) for processing our RNA representation. To handle the different modalities of , our model consists of a base network that feeds into 3 headers for each data component: sequence, coordinates and 3D features. The sequence header predicts a probability vector , while coordinate and feature headers predict equivariant variables and . We train the model to reconstruct the original structure and sequence :
@@ 3 Results
We leverage the RNASolo dataset (Adamczyk et al., 2022) which consists of extracted individual RNA structures from the Protein Databank (PDB) (Berman et al., 2000) to train our model. Following (Anand et al., 2024), we filter the full dataset to resolution and sequence lengths between 40 and 150 to a total dataset size of 6090 data points. The dataset exhibits significant length imbalances which led to biased model performance initially, to address this, we implemented a length-balanced sampling that ensures uniform representation of RNA sequences for training as described in Appendix A.1. We train our model with batch size of 64 for 120k steps on 4 GPUs. Our model’s flow process is trained on 300 timesteps. To evaluate our model, we follow (Anand et al., 2024) and sample 50 RNA structures at each sequence length from 40-150 with step size of 10.
@@ 3.1 Chemical Validity
To assess the chemical validity of our generated structures, we analyzed key geometric parameters that define RNA backbone and base conformations. Using MDAnalysis (Gowers et al., 2016) (Michaud-Agrawal et al., 2011) we computed all dihedral angles and the pseudo-angle for the ribose pucker. Figure 3 shows the distributions of the dihedral angles across the training data (representing experimentally determined structures) and 50 randomly selected RiboGen structures, 5 from each sequence length. The results demonstrate that for most angles RiboGen generated structures capture the dihedral angles distributions and the general trends of the training set, though with some discrepancies. Our RiboGen structures show broader distribution across the Alpha angle indicating mild divergence from the experimental data in this specific torsion. Overall these results suggest that RiboGen has successfully learned the geometric constraints of RNA molecules.
@@ 3.2 Self-Consistency
To evaluate the quality and biological plausibility of our generated RNA, we employed a self consistency validation process. For each generated RNA molecule, we extracted its sequence and used Boltz-1 (Wohlwend et al., 2024) to obtain a reference structure. We quantify structural similarity between our generated structure and Boltz’s structure using two complementary metrics: Root Mean Square Deviation (RMSD) and Template Modeling score (TM-score). We calculated these metrics across all 50 samples for each sequence length. Following (Anand et al., 2024), in Figure 5(a) and (b) we report results on 10 best-performing samples per length. Our predictions scTM scores in 5(b) demonstrate lower variance than in Anand et al. (2024), suggesting better consistency and generalization across different lengths. In Figure 5(c) we observe that RiboGen achieves higher median TM-scores than RNA-FrameFlow for most sequence lengths above 70 nucleotides. This may highlight the benefits of co-generation, especially for longer RNA.
@@ 3.3 Structural Evaluation
To evaluate RiboGen’s performance in generating valid RNA structure-sequence pairs, we utilized the metrics from the evaluation suite proposed by (Anand et al., 2024). While RNA-FrameFlow focuses on structure generation alone, RiboGen jointly generates both the sequence and its corresponding structure. Therefore, instead of using gRNAde (Joshi et al., 2025) sequences and corresponding RhoFold (Shen et al., 2024) structures to calculate TM-score, we folded the co-generated sequences using Boltz-1 and computed the TM-score after aligning them to the generated structures. While RNA-FrameFlow employs an additional 8-shot inverse folding process when calling gRNAde eight times for each backbone, RiboGen performs one-shot co-generation of both sequence and structure in a single sampling process. This highlights RiboGen’s potential for simpler and more efficient RNA design workflows, avoiding the need for expensive post-hoc sequence inference. Following (Anand et al., 2024), samples with were considered valid. To assess diversity, we measured the number of unique qTM clusters among the valid samples and normalized this by the total number of valid samples. Although RiboGen is sampling jointly RNA sequences and structures, it achieved performance on par with RNA-FrameFlow in terms of backbone validity and diversity. Our results demonstrate competitive metrics and efficient sampling, validating RiboGen as a promising baseline for joint RNA structure-sequence generation.
Model | Sampling Steps | % Validity | Diversity | Time (s) |
RiboGen | 100 | 27.17 | 0.604 | 1.18 |
200 | 32.17 | 0.553 | 4.50 | |
300 | 34.17 | 0.585 | 9.06 | |
RNA-FrameFlow * | 10 | 16.7 | 0.62 | – |
50 | 41.0 | 0.61 | 4.74 | |
100 | 20.0 | 0.61 | – | |
MMDiff * | 100 | 0 | – | 27.30 |
@@ Acknowledgments
This research was made possible through the support of the Eleven Eleven Foundation, the Center for Bits and Atoms, and the MIT Media Lab Consortium. Their support was fundamental in enabling this work.

### v5

@@ 1 Introduction
Ribonucleic acid (RNA) is a fundamental biomolecule that stands at the intersection of modern biology and the origins of life. RNA has proven to be a versatile molecule, playing key roles in messaging (crick1970centraldogma), catalytic functions (guerrier1983rna), regulation, and diverse biological processes through its complex 3D structures (fire1998rna). While traditional computational methods faced limitations in decoding RNA structure and facilitating RNA design, deep learning emerged as powerful approach to accurately predict RNA structures, enhance RNA engineering, and unlock new insights into its functional roles. Existing deep learning models often predict RNA structure from sequence or design sequences for target structures independently. However, the ability to simultaneously generate both sequence and structure remains a largely unexplored space in deep learning-based modeling of RNA. This co-generation capability can enable exploration of the sequence-structure landscape in novel ways. To address this gap, in this paper we introduce RiboGen for joint generation of all-atom structure and sequence of RNA. RiboGen is based on a Multiflow (campbell2024generative) model that draws on Flow Matching (lipman2022flow; liu2022flow) and Discrete Flow (gat2024discreteflowmatching; campbell2024generative) for its generation. We train a large model and evaluate its chemical validity and self-consistency. Our results showcase the capabilities of RiboGen for generation and highlight that co-generation models are a promising avenue for modeling of RNA.
@@ 1.1 Related Work
Previous work in deep learning for RNA design has made significant progress in computational approaches to predict RNA structures (shen2024accurate; abramson2024), RNA-RNA interactions, and the design of novel RNA sequences. Recent progress in generative modeling for RNA focuses on sequence and structure generation through Denoising Diffusion Probabilistic Models (DDPM) (ho2020denoising) or Flow Matching (lipman2022flow). MMDiff (morehead2023jointsequencestructuregenerationnucleic) uses discrete DDPM to co-generate sequence and structure of RNA, DNA and proteins. Our approach instead employs Flow Matching and its discrete variant (campbell2024generative). RNA-FrameFlow (anand2024rna) represents RNA through rigid body frames and uses flow matching to generate 3D backbones, employing inverse folding model gRNAde (joshi2023grnade) to obtain sequences. In contrast, while similarly using Flow Matching for 3D generation, our approach additionally models the discrete sequence components of RNA generation. RNAFlow nori2024rnaflow uses a GNN conditioned on protein structure and sequence to generate RNA sequences, which are then processed by RoseTTAFold2NA Baek2024 to predict backbone structure; this method additionally conditions on protein structure as input. Our approach instead focuses on isolated RNA, learning unconditional direct sequence-structure generation. Recent application of the Multiflow (campbell2024generative) framework to protein sequence-structure design has demonstrated the power of joint generation. Our approach builds upon on these insights from protein design, and adapts them to the RNA design domain by enabling the joint generation of RNA sequences and all-atom structures.
@@ 2.2 Flow Matching
To model the distribution of 3D RNA coordinates and features we use Flow Matching (lipman2022flow; liu2022flow; albergo2023stochastic). Flow Matching parameterizes a conditional probability path on time by learning a conditional velocity field that transforms samples from a prior distribution to a target data distribution . To learn this transport, we use the standard form of Flow Matching to obtain a noised version of via the linear interpolant and its associated velocity:
We build our model to reconstruct the target from its noised counterpart . We follow the reparameterization of (jing2024alphafold; pooladian2023multisampleflowmatchingstraightening) and obtain the learned conditional velocity through:
@@ 2.3 Discrete Flow Matching
While the standard form of Flow Matching is effective for continuous data, it is not appropriate for categorical domains. Hence, for modeling the RNA sequence we employ the extended framework of Discrete Flow Matching (gat2024discreteflowmatching; campbell2024generative). In this setting, the sequence data is described over a vocabulary . We parameterize a discrete flow by describing the velocity field over a probability vector on this categorical space:
@@ 2.4 Multiflow
To generate full RNA representations, we use Multiflow (campbell2024generative) and train a neural network to learn the multimodal velocity field given jointly noised data and time (Figure 1.b). We employ the standard Flow Matching for and and its discrete counterpart for . This decomposition enables the model to separately capture the distributions of sequence, backbone, and atomic positions, allowing for unconditional generation or conditional generation based on specific structural or sequence constraints (Figure 2.a), such as structure prediction or inverse folding.
@@ 2.5 Architecture and Training
We use Euclidean-Equivariant Neural Networks (geiger2022e3nn) for processing our RNA representation. To handle the different modalities of , our model consists of a base network that feeds into 3 headers for each data component: sequence, coordinates and 3D features. The sequence header predicts a probability vector , while coordinate and feature headers predict equivariant variables and . We train the model to reconstruct the original structure and sequence :
@@ 3 Results
We leverage the RNASolo dataset (adamczyk2022rnasolo) which consists of extracted individual RNA structures from the Protein Databank (PDB) (berman2000protein) to train our model. Following (anand2024rna), we filter the full dataset to resolution and sequence lengths between 40 and 150 to a total dataset size of 6090 data points. The dataset exhibits significant length imbalances which led to biased model performance initially, to address this, we implemented a length-balanced sampling that ensures uniform representation of RNA sequences for training as described in Appendix A.1. We train our model with batch size of 64 for 120k steps on 4 GPUs. Our model’s flow process is trained on 300 timesteps. To evaluate our model, we follow (anand2024rna) and sample 50 RNA structures at each sequence length from 40-150 with step size of 10.
@@ 3.1 Chemical Validity
To assess the chemical validity of our generated structures, we analyzed key geometric parameters that define RNA backbone and base conformations. Using MDAnalysis (Gowers2016) (Michaud-Agrawal2011) we computed all dihedral angles and the pseudo-angle for the ribose pucker. Figure 3 shows the distributions of the dihedral angles across the training data (representing experimentally determined structures) and 50 randomly selected RiboGen structures, 5 from each sequence length. The results demonstrate that for most angles RiboGen generated structures capture the dihedral angles distributions and the general trends of the training set, though with some discrepancies. Our RiboGen structures show broader distribution across the Alpha angle indicating mild divergence from the experimental data in this specific torsion. Overall these results suggest that RiboGen has successfully learned the geometric constraints of RNA molecules.
@@ 3.2 Self-Consistency
To evaluate the quality and biological plausibility of our generated RNA, we employed a self consistency validation process. For each generated RNA molecule, we extracted its sequence and used Boltz-1 (boltz1) to obtain a reference structure. We quantify structural similarity between our generated structure and Boltz’s structure using two complementary metrics: Root Mean Square Deviation (RMSD) and Template Modeling score (TM-score). We calculated these metrics across all 50 samples for each sequence length. Following (anand2024rna), in Figure 5(a) and (b) we report results on 10 best-performing samples per length. Our predictions scTM scores in 5(b) demonstrate lower variance than in anand2024rna, suggesting better consistency and generalization across different lengths. In Figure 5(c) we observe that RiboGen achieves higher median TM-scores than RNA-FrameFlow for most sequence lengths above 70 nucleotides. This may highlight the benefits of co-generation, especially for longer RNA.
@@ 3.3 Structural Evaluation
To evaluate RiboGen’s performance in generating valid RNA structure-sequence pairs, we utilized the metrics from the evaluation suite proposed by (anand2024rna). While RNA-FrameFlow focuses on structure generation alone, RiboGen jointly generates both the sequence and its corresponding structure. Therefore, instead of using gRNAde (joshi2025grnadegeometricdeeplearning) sequences and corresponding RhoFold (shen2024accurate) structures to calculate TM-score, we folded the co-generated sequences using Boltz-1 and computed the TM-score after aligning them to the generated structures. While RNA-FrameFlow employs an additional 8-shot inverse folding process when calling gRNAde eight times for each backbone, RiboGen performs one-shot co-generation of both sequence and structure in a single sampling process. This highlights RiboGen’s potential for simpler and more efficient RNA design workflows, avoiding the need for expensive post-hoc sequence inference. Following (anand2024rna), samples with were considered valid. To assess diversity, we measured the number of unique qTM clusters among the valid samples and normalized this by the total number of valid samples. Although RiboGen is sampling jointly RNA sequences and structures, it achieved performance on par with RNA-FrameFlow in terms of backbone validity. Our results demonstrate competitive metrics and efficient sampling, validating RiboGen as a promising baseline for joint RNA structure-sequence generation.
| Model | Sampling Steps | % Validity | Time (s) |
| RiboGen | 100 | 27.17 | 1.18 |
| 200 | 32.17 | 4.50 | |
| 300 | 34.17 | 9.06 | |
| RNA-FrameFlow * | 10 | 16.7 | – |
| 50 | 41.0 | 4.74 | |
| 100 | 20.0 | – | |
| MMDiff * | 100 | 0 | 27.30 |
@@ Acknowledgments
This research was made possible through the support of the Eleven Eleven Foundation, the Center for Bits and Atoms, and the MIT Media Lab Consortium